# VIEWER: an extensible visual analytics framework for enhancing mental healthcare

### v1

@@ Abstract
Objective: To design and implement VIEWER, a versatile toolkit for visual analytics of clinical data, and to systematically evaluate its effectiveness across various clinical applications while gathering feedback for iterative improvements.
Materials and Methods: VIEWER is an open-source and extensible toolkit that employs distributed natural language processing and interactive visualisation techniques to facilitate the rapid design, development, and deployment of clinical information retrieval, analysis, and visualisation at the point of care. Through an iterative and collaborative participatory design approach, VIEWER was designed and implemented in a large mental health institution, where its clinical utility and effectiveness were assessed using both quantitative and qualitative methods.
Results: VIEWER provides interactive, problem-focused, and comprehensive views of longitudinal patient data from a combination of structured clinical data and unstructured clinical notes. Despite a relatively short adoption period and users’ initial unfamiliarity, VIEWER significantly improved performance and task completion speed compared to the standard clinical information system. Users and stakeholders reported high satisfaction and expressed strong interest in incorporating VIEWER into their daily practice.
@@
Introduction
As the volume of information in patient records continues to grow, clinicians face an overwhelming amount of complex raw data, which can exceed the capacity of human cognition to process without error [1]. This issue of information overload has been widely noted across various care settings, including both primary and secondary care, which may result in diagnostic or treatment errors [2, 3]. Moreover, the growing demand for health information exchange among healthcare providers, and between providers and stakeholders such as public health agencies and regulators, has made patient data review and administrative reporting increasingly complex and time-consuming [4]. This is particularly evident in mental healthcare, where patients often require ongoing mental health support and coordinated care for physical comorbidities and social support [5, 6]. As a result, clinicians are forced to spend valuable time assembling disparate data points to create a coherent view for a patient’s care, which can lead to inefficiency and delayed care [7].
@@ Introduction
However, existing research has mainly focused on physical healthcare settings, with limited attention given to mental healthcare [23]. Compared to physical health, visual analysis of mental health data presents greater complexity for two key reasons. First, mental illnesses often encompass both medical, psychological and social dimensions [24], requiring a broader range of data to achieve a comprehensive assessment of an individual’s mental health. This highlights the need of an extensible toolkit that can facilitate the rapid design, development, and implementation of visual analytics across diverse data sources and types in mental healthcare. Second, treating mental illnesses typically involves a combination of interventions, including medication, psychotherapy, lifestyle changes, and social support services [6]. Unlike physical conditions, where structured data (such as blood assays and other characterizations) is more salient, information about mental health presentations, relevant contextual factors, interventions, and outcomes is often documented in unstructured text (e.g., clinical notes and correspondences) [25]. Integrating this unstructured information into visual analytics pipelines is more challenging compared to structured data such as numerical or categorical variables.
@@
Another notable limitation of many previous studies is their generalisability. First, most studies have focused on a specific clinical task [7, 16] or disease [18, 26]. However, in real-world clinical settings, patients rarely present with a single disease or single risk factor that may affect their health [27]. There is a lack of integrated visual-analytics solutions that can systematically address the varied challenges faced by different healthcare stakeholders, including clinicians and patients [22]. Second, existing solutions are often developed using proprietary or custom tools internally developed by hospitals or EHR system providers [7, 26], leading to limited interoperability with other tools and restricted applicability in different settings [8]. Finally, for visual-analytics solutions to achieve better outcomes and effective adoption, they should be user-friendly, require minimal training, offer evidence-based recommendations, and integrate smoothly into clinicians’ workflows [27]. Despite extensive focus on technical development in previous studies, there has been little effort to systematically develop and evaluate these solutions within the complex workflows of clinicians, nurses, and other health workers while considering their varied clinical priorities and patient needs [28].
@@
In this work, we present VIEWER (Visual & Interactive Engagement With Electronic Records), an open-source, cost-effective, and extensible toolkit created for the rapid design, development, and deployment of clinical information retrieval, analysis, and visualisation for supporting clinical decision making. VIEWER is an EHR-agnostic framework that utilises distributed information extraction pipelines, leveraging natural language processing (NLP) methods and open-source visualisation techniques to enable comprehensive search and visual analytics of a comprehensive health record from both structured and unstructured patient data within a health institution, rather than a curated dataset for a specific disease or patient cohort. We also systematically describe our interdisciplinary and collaborative approach to participatory design where we designed, implemented and evaluated VIEWER within one of the largest National Health Services (NHS) Trusts for mental health in the United Kingdom (UK), aimed at addressing the diverse challenges faced by multiple healthcare stakeholders. The stakeholders included: (i) clinicians who need to synthesise disparate data to understand a patient’s condition within the context of their medical history, (ii) managers who require data-driven insights to optimise resource allocation and identify unmet needs, (iii) researchers who seek to understand disparities in outcomes across populations, and (iv) patients who wish to utilise their own medical data for self-monitoring. Our evaluation demonstrates the effectiveness of VIEWER in enhancing patient care in real-world clinical use cases and provides valuable insights into working collaboratively with clinical workers, researchers, patients and carers, and informaticians to iteratively refine and optimise informatics solutions for improved patient care.
@@
This work presents VIEWER, an extensible visual-analytics system for improving clinical data accessibility and presentation for clinical decision support. VIEWER leverages distributed NLP-based information retrieval and extraction modules, and open-source visualisation services to provide visually interactive, problem-oriented, and holistic presentations of the patient record. In this work, a participatory, iterative, and interdisciplinary collaborative design approach was adopted to develop an initial prototype, implement it to extend the EHR ecosystem at the South London and Maudsley NHS Foundation Trust (SLaM), and continuously refine the ecosystem based on output from user feedback cycles. Evaluations were also conducted to assess VIEWER’s utility and effectiveness in supporting healthcare professionals across various clinical workflows, such as patient health checks, medication reviews, and caseload management, thereby providing opportunities to identify areas for improvement of the implementation in its future iterations.
Materials and Methods
Setting
The VIEWER platform was developed within the National Institute for Health and Care Research (NIHR) Maudsley Biomedical Research Centre (BRC) and has been deployed at SLaM, the UK’s largest mental health NHS Trust. SLaM provides more than 240 secondary mental health services for a local population of 1.3 million residents in South London. In addition, SLaM provides more than 50 national and specialist services for people across the UK and beyond. Each year, the Trust provides inpatient care for over 5,000 people and treat more than 40,000 patients.
@@
Like other EHR systems [7], ePJS is a Web-based application designed for entering and reviewing patient information, with a specific focus on displaying plain textual data rather than visual graphics. Patient summaries within a clinician’s caseload are presented in a tabular format, where each row provides a link to a patient’s profile. This profile contains various tabs corresponding to different components of patient details, including clinical notes, laboratory tests, diagnoses, and medication prescriptions. All data from the clinical record are fine-grained information at the individual patient level, with no aggregation of temporal information for the same patient or aggregation at the team or population levels. A basic search function allows users to search based on document type, test type, recording date, patient name, and their unique identifiers such as the national NHS number or local Trust identifier (ID). However, the clinical data in free-text documents, constituting the majority of the EHR, are less amenable for search queries and less accessible to clinicians.
@@
Table 1: Needs identified for different stakeholders.
@@
Design approach
VIEWER was initiated by an interdisciplinary team, which consisted of two clinicians, two computer scientists, and two health informaticians, at SLaM in January 2020, aiming to enhance data-driven clinical decision support and promote population health management within the Trust. To ensure that the perspectives, needs, and preferences of all stakeholders are considered, an iterative participatory design approach [31, 32], a widely used approach for clinical decision support systems [33, 31], was adopted throughout the design, development, and deployment stages. This approach involves a collaborative process, whereby all stakeholders engage with each other to co-create and co-improve an artefact uring development and deployment, with the aim of developing solutions that are usable, useful, and effective.
Our design process included three phases. In phase 1, stakeholder needs were identified and discussed. Table 1 summarises the needs identified from different categories of stakeholders. These needs stem from 1) brainstorming sessions within the development team, and 2) feedback obtained from prototype demo sessions and design workshops involving the various stakeholders, including clinical workers. Phase 2 involved the design and development of a series of prototypes, leveraging interdisciplinary skills from the development team. The team held routine weekly meetings to identify and prioritise areas for improvement, brainstorm practical solutions, and plan the next iteration of development. Each prototype underwent testing and was eventually migrated into a production environment after reaching consensus by all relevant stakeholders. In phase 3, each new prototype was piloted and evaluated in a real clinical setting. During this phase, demos and training materials, including user guides, screenshots, and recorded video tutorials, were provided to pilot teams. Feedback was actively collected during demos, training sessions, and feedback channels (e.g. emails and review sessions) to guide the next iteration of development.
@@
Overall architecture
Figure 1 illustrates the architecture and pipelines used in VIEWER. Patient data, comprising both structured and unstructured information, is initially collected and recorded in a SQL database using ePJS, SLaM’s EHR system, at the point of care. The updated EHR in ePJS is synchronised daily with a staging SQL database in the CRIS system using Extract, Transform, and Load (ETL) scripts. This staging environment allows the ETL processes to not only handle raw data in ePJS, but also to integrate NLP pipelines for extracting key medical information (e.g. symptoms, interventions, outcomes, and contextual factors) from free-text notes [34] [34], and data linkage connecting local data with other clinical data sources to enhance the richness of EHR [35]. We designed and tested the ETL pipelines based on de-identifiable data in CRIS and plug in these pipelines to raw EHR from ePJS to enable clinical use of data.
@@
Figure 2: VIEWER data models. The physical data mode illustrates physical data stores of the source EHR data, the conceptual data model describes the conceptual structure of entity classes and the logical data model describes the semantic and logical relationships among entities.
Data model
To meet the varying needs of different stakeholders, we have crafted a comprehensive data model, encompassing a diverse array of patient attributes, clinical activities, and outcomes. As shown in Figure 2, we first identify relevant data sources within the source EHR, including both structured information and unstructured text data. These sources are then transformed and conceptualised into entity classes based on local settings and project-specific requirements.
@@
Data security and audit
@@
Results
Patient characteristics
As of a recent census date on March 25th 2024, VIEWER contains information for all patients (n=409,870) who had a team episode or ward stay at SLaM, where 51,057 were active caseload, namely those currently managed by the Trust. Table 2 summarises patient characteristics for active and inactive/discharged caseload. Compared to the inactive caseload, patients in the active caseload tend to younger and have a higher proportion from the Black and Mixed ethnicity groups. Also, active caseload has a higher prevalence of severe mental illnesses, such as schizophrenia (ICD 10 code F2), bipolar disorder (F31) and major depressive disorder, and a higher proportion of living in the SLaM catchment area.
@@ Results
Physical health monitoring
@@
The first use case of VIEWER was to support annual physical health check in mental healthcare. Compared to the general population, people with mental illnesses often face poorer physical health [37], which has contributed to reduced quality of life and a shortened life expectancy by 10-20 years for people with severe mental illnesses [38, 39]. To enhance overall health and decrease premature mortality among this group, clinical guidelines recommend annual physical health checks to promptly identify and manage physical health conditions [^3] https://www.england.nhs.uk/long-read/improving-the-physical-health-of-people-living-with-severe-mental-illness/. These assessments typically include the evaluation of multiple parameters, including blood pressure, body mass index (BMI), blood glucose or HbA1c levels, lipid profile, smoking status, alcohol consumption, and lifestyle factors like diet and physical activity levels, which can be time consuming.
@@
VIEWER has served as the primary system supporting a team of six members in overseeing physical health checks for the entire active caseload of over 36,000 patients within the Trust. VIEWER provided a comprehensive overview of physical health checks by leveraging NLP-based information extraction methods. Through interactive visualisations, team members can readily identify patients who require a physical health assessment. Each identified patient is provided with a hyperlink that directs clinicians to their profile within the EHR system, facilitating actions such as scheduling a check or contacting GPs for additional information. After completion, these checks are recorded in the source EHR system and then incorporated into VIEWER. This systematic approach has streamlined the check process and resulted in an increase in completed assessments. For example, Figure 4 shows the rise in the number of measures completed by the team after adopting VIEWER as a systematic approach within a short period of two months.
Medication review

### v2

@@ Abstract
Objective: A proof-of-concept study aimed at designing and implementing VIEWER, a versatile toolkit for visual analytics of clinical data, and systematically evaluating its effectiveness across various clinical applications while gathering feedback for iterative improvements.
Materials and Methods: VIEWER is an open-source and extensible toolkit that employs natural language processing and interactive visualisation techniques to facilitate the rapid design, development, and deployment of clinical information retrieval, analysis, and visualisation at the point of care. Through an iterative and collaborative participatory design approach, VIEWER was designed and implemented in one of the UK’s largest NHS mental health Trusts, where its clinical utility and effectiveness were assessed using both quantitative and qualitative methods.
Results: VIEWER provides interactive, problem-focused, and comprehensive views of longitudinal patient data (n=409,870) from a combination of structured clinical data and unstructured clinical notes. Despite a relatively short adoption period and users’ initial unfamiliarity, VIEWER significantly improved performance and task completion speed compared to the standard clinical information system. More than 1,000 users and partners in the hospital tested and used VIEWER, reporting high satisfaction and expressed strong interest in incorporating VIEWER into their daily practice.
@@
1 Introduction
As the volume of information in patient records continues to grow, clinicians face an overwhelming amount of complex raw data, which can exceed the capacity of human cognition to process without error [1]. This issue of information overload has been widely noted across various care settings, including both primary and secondary care, which may result in diagnostic or treatment errors [2, 3]. Moreover, the growing demand for health information exchange among healthcare providers, and between providers and partners such as public health agencies and regulators, has made patient data review and administrative reporting increasingly complex and time-consuming [4]. This is particularly evident in mental healthcare, where patients often require ongoing mental health support and coordinated care for physical comorbidities and social support [5, 6]. As a result, clinicians—including psychiatrists, psychologists, pharmacists, nurses, and professionals offering occupational and social support in mental health services—are forced to spend valuable time assembling disparate data points to create a coherent view for a patient’s care, which can lead to inefficiency and delayed care [7].
@@ 1 Introduction
However, existing research has mainly focused on physical healthcare settings, with limited attention given to mental healthcare [23]. Compared to physical health, visual analysis of mental health data presents greater complexity for two key reasons. First, mental health conditions often encompass both medical, psychological and social dimensions [24], requiring a broader range of data to achieve a comprehensive assessment of an individual’s mental health. This highlights the need of an extensible toolkit that can facilitate the rapid design, development, and implementation of visual analytics across diverse data sources and types in mental healthcare. Second, treating mental health conditions typically involves a combination of interventions, including medication, psychotherapy, lifestyle changes, and social support services [6]. Unlike physical conditions, where structured data (such as blood assays and other characterizations) is more salient, mental health assessments also rely on vast patient-reported information as quantitative measures to evaluate various aspects of mental health, including mental health presentations, relevant contextual factors, interventions, and outcomes [25]. Such information is often documented in unstructured text, such as clinical notes and correspondences [26]. Integrating this unstructured information into visual analytics pipelines is more challenging compared to structured data such as numerical or categorical variables. Thus, tools developed for physical health management in previous studies may not effectively process and represent mental healthcare data.
@@
Another notable limitation of many previous studies is their generalisability. First, most studies have focused on a specific clinical task [7, 16] or disease [18, 27]. However, in numerous clinical settings, patients rarely present with a single disease or single risk factor that may affect their health [28]. There is a lack of integrated visual-analytics solutions that can systematically address the varied challenges faced by different healthcare partners, including clinicians and patients [22]. Second, existing solutions are often developed using proprietary or custom tools internally developed by hospitals or EHR system providers [7, 27], leading to limited interoperability with other tools and restricted applicability in different settings [8]. Finally, for visual-analytics solutions to achieve better outcomes and effective adoption, they should be user-friendly, require minimal training, offer evidence-based recommendations, and integrate smoothly into clinicians’ workflows [28]. Despite extensive focus on technical development in previous studies, there has been little effort to systematically develop and evaluate these solutions within the complex workflows of clinicians, nurses, and other health workers while considering their varied clinical priorities and patient needs [29].
@@
In this work, we present VIEWER (Visual & Interactive Engagement With Electronic Records), an open-source, cost-effective, and extensible toolkit created for the rapid design, development, and deployment of clinical information retrieval, analysis, and visualisation for supporting clinical decision making. VIEWER is an EHR-agnostic framework that utilises distributed information extraction pipelines, leveraging natural language processing (NLP) methods and open-source visualisation techniques to enable comprehensive search and visual analytics of a comprehensive health record from both structured and unstructured patient data within a health institution, rather than a curated dataset for a specific disease or patient cohort. We also systematically describe our interdisciplinary and collaborative approach to participatory design where we designed, implemented and evaluated VIEWER on top of a bespoke EHR system within one of the largest National Health Services (NHS) Trusts for mental health in the United Kingdom (UK), aimed at addressing the diverse challenges faced by multiple healthcare partners. The partners included: (i) clinicians who need to synthesise disparate data to understand a patient’s condition within the context of their medical history, (ii) managers who require data-driven insights to optimise resource allocation and identify unmet needs, (iii) researchers who seek to understand disparities in outcomes across populations, and (iv) patients who wish to utilise their own medical data for self-monitoring. Our evaluation demonstrates the effectiveness of VIEWER in enhancing patient care in real-world clinical use cases and provides valuable insights into working collaboratively with clinical workers, researchers, patients and carers, and informaticians to iteratively refine and optimise informatics solutions for improved patient care.
@@
This paper describes the development and implementation of VIEWER, an extensible visual-analytics system for improving clinical data accessibility and presentation for clinical decision support, and the system’s proof-of-concept applications within mental health services. We first describe a participatory, iterative, and interdisciplinary collaborative design approach used to develop an initial prototype of VIEWER, implement it to extend the EHR ecosystem at the South London and Maudsley NHS Foundation Trust (SLaM), and continuously refine the ecosystem based on output from user feedback cycles. We then detail the technical components of VIEWER, which includes distributed NLP-based information retrieval and extraction modules, and open-source visualisation services to provide visually interactive, problem-oriented, and holistic presentations of the patient record. Finally, we conduct evaluations to assess VIEWER’s utility and effectiveness in supporting healthcare professionals across various clinical workflows, such as patient health checks, medication reviews, and caseload management, thereby providing opportunities to identify areas for improvement of the implementation in its future iterations.
2 Materials and Methods
2.1 Setting
The VIEWER platform was developed within the National Institute for Health and Care Research (NIHR) Maudsley Biomedical Research Centre (BRC) and has been deployed at SLaM, the UK’s largest mental health NHS Trust. SLaM provides more than 240 secondary mental health services [^1] In the UK, secondary mental health services refer to mental health services in secondary care, typically including hospitals, some psychological wellbeing services, community mental health teams (CMHTs), crisis resolution and home treatment teams (CRHTs), assertive outreach teams and early intervention teams. for a local population of 1.3 million residents in South London. In addition, SLaM provides more than 50 national and specialist services for people across the UK and beyond. Each year, the Trust provides inpatient care for over 5,000 people and treat more than 40,000 patients.
@@ 2.1 Setting
Like other EHR systems [7], ePJS is a Web-based application designed for entering and reviewing patient information, with a specific focus on displaying plain textual data rather than visual graphics. Patient summaries within a clinician’s caseload are presented in a tabular format, where each row provides a link to a patient’s profile. This profile contains various tabs corresponding to different components of patient details, including clinical notes, laboratory tests, diagnoses, and medication prescriptions. All data from the clinical record are fine-grained information at the individual patient level, with no aggregation at the team/population levels or aggregation of temporal information for the same patient. A basic search function allows users to search based on document type, test type, recording date, patient name, and their unique identifiers such as the national NHS number or local Trust identifier (patient IDs). However, the clinical data in free-text documents, constituting the majority of the EHR, are less amenable for search queries and less accessible to clinicians.
@@
Table 1: Needs identified for different partners.
@@
2.2 Interdisciplinary collaboration
VIEWER was initiated by an interdisciplinary team, which consisted of two clinicians, two computer scientists, and two health informaticians, at SLaM in January 2020, aiming to enhance data-driven clinical decision support and promote population health management within the Trust. To ensure that the perspectives, needs, and preferences of all partners are considered, an iterative participatory design approach [32, 33], a widely used approach for clinical decision support systems [34, 32], was adopted throughout the design, development, and deployment stages. This approach involves a collaborative process, whereby all partners engage with each other to co-create and co-improve the development and deployment, with the aim of developing solutions that are usable, useful, and effective.
Our design process included three phases. In phase 1, partner needs were identified and discussed. These needs stem from 1) brainstorming sessions within the development team, and 2) feedback obtained from prototype demo sessions and design workshops involving the various partners, including clinical workers. Table 1 summarises the resulting needs identified from different categories of partners in this phase. Phase 2 involved the design and development of a series of prototypes, leveraging interdisciplinary skills from the development team. The team held routine weekly meetings to identify and prioritise areas for improvement, brainstorm practical solutions, and plan the next iteration of development. Each prototype underwent testing, which involves: 1) validating the accuracy of information presented in VIEWER by comparing it with the raw data in ePJS, particularly for data processed by NLP pipelines such as medication extraction [35], 2) assessing the appropriateness of the information’s visualization, including the suitability of visual graph types and labeling, and 3) the accessibility of the visual interface, focusing on logical design, layout coherence, and integration with other system components. A prototype was eventually migrated into a production environment after reaching consensus by all relevant partners. Key results from this phase are showcased through the user interfaces detailed in Section 3.3. In phase 3, each new prototype was piloted and evaluated in a real clinical setting. During this phase, demos and training materials, including user guides, screenshots, and recorded video tutorials, were provided to pilot teams. The pilot sessions were centered on specific use cases and involved participants with diverse backgrounds, demographics, and areas of expertise. Participation was voluntary from teams that requested involvement or expressed interest in a given use case. No formal selection process was implemented, ensuring organic representation to promote diversity, inclusivity, and equity in engagement. Feedback was actively collected during demos, training sessions, and feedback channels (e.g. emails and review sessions) to guide the next iteration of development. The key outcomes from this phase are outlined in the use cases in Section 3.4 and the system usability findings in Section 3.5.
It is important to highlight that the design process followed an iterative approach. Requests from partners and feedback from end-users were continuously gathered and prioritized by the development team to refine a current version and identify areas for improvement in subsequent versions. This process facilitated ongoing feedback loops, which ultimately led to the creation of a final design that more closely aligned with the users’ needs and expectations.
@@ 2.2 Interdisciplinary collaboration
2.3 Overall architecture
Figure 1 illustrates the architecture and pipelines used in VIEWER. Patient data, comprising both structured and unstructured information, is initially collected and recorded in a SQL database using ePJS, SLaM’s EHR system, at the point of care. The updated EHR in ePJS is synchronised daily with a staging SQL database in the CRIS system using Extract, Transform, and Load (ETL) scripts. This staging environment allows the ETL processes to not only handle raw data in ePJS, but also to integrate NLP pipelines for extracting key medical information (e.g. symptoms, interventions, outcomes, and contextual factors) from free-text notes [36], and data linkage connecting local data with other clinical data sources to enhance the richness of EHR [37]. The NLP pipelines extract mentions of medical entities, such as symptoms and interventions, from clinical notes using rule-based and machine learning approaches built on the General Architecture for Text Engineering (GATE) software. GATE is a comprehensive suite of tools used for diverse NLP tasks such as text parser, morphology, tagging and information extraction [38]. These NLP pipelines have been regularly validated (with precision and recall metrics reported and updated in [36]) and deployed to extract clinical information from routinely collected EHR in CRIS. Details on the development and validation of the NLP applications within CRIS are documented in previous publications [26, 39, 38] and online resources [36]. We designed and tested the ETL pipelines based on de-identifiable data in CRIS and plug in these pipelines to raw EHR from ePJS to enable clinical use of data.
@@ 2.3 Overall architecture
Figure 2: VIEWER data models. The physical data model illustrates physical data stores of the source EHR data, the conceptual data model describes the conceptual structure of entity classes and the logical data model describes the semantic and logical relationships among entities.
2.4 Data model
To meet the varying needs of different partners, we have crafted a comprehensive data model, encompassing a diverse array of patient attributes, clinical activities, and outcomes. As shown in Figure 2, we first identify relevant data sources within the source EHR, including both structured information and unstructured text data. These sources are then transformed and conceptualised into entity classes based on local settings and project-specific requirements.
@@ 2.5 Interface design
  • Learnability and error recovery: To enhance learnability, employ consistent design elements such as icons, buttons, terminology, and structured layouts, while avoiding technical jargon. Additionally, incorporate default interfaces with “undo” options and accessible help guides to facilitate error recovery and ensure users can easily correct mistakes.
@@ 2.6 Scalability
2.7 Data security and audit
@@
3 Results
3.1 Patient characteristics
As of a recent census date on March 25th 2024, VIEWER ingested information for all patients (n=409,870) who had a team episode or ward stay at SLaM, where 51,057 were active caseload, namely those currently managed by the Trust. Table 2 summarises the patient characteristics for both active and inactive/discharged caseloads that have been integrated into VIEWER. Compared to the inactive caseload, patients in the active caseload tend to younger and have a higher proportion from the Black and Mixed ethnicity groups. Also, active caseload has a higher prevalence of severe mental health conditions, such as schizophrenia (ICD 10 code F2), bipolar disorder (F31) and major depressive disorder, and a higher proportion of living in the SLaM catchment area.
@@ 3.4 Clinical applications
3.4.1 Physical health monitoring
@@
The first use case of VIEWER was to support annual physical health check for all active patients within the Trust, particularly for those living with severe mental health conditions or undergoing long-term psychiatric medication treatment. Compared to the general population, people with mental health conditions often face poorer physical health [41], which has contributed to reduced quality of life and a shortened life expectancy by 10-20 years for people with severe mental health conditions [42, 43]. This disparity is partly due to the potential side effects of psychiatric medications, such as weight gain, metabolic syndrome, cardiovascular risks, and other chronic conditions [44, 45]. To enhance overall health and decrease premature mortality among this group, clinical guidelines recommend annual physical health checks to promptly identify and manage physical health conditions [^4] https://www.england.nhs.uk/long-read/improving-the-physical-health-of-people-living-with-severe-mental-illness/. These assessments typically include the evaluation of multiple parameters, including blood pressure, body mass index (BMI), blood glucose or HbA1c levels, lipid profile, smoking status, alcohol consumption, and lifestyle factors like diet and physical activity levels, which can be time consuming.
@@
VIEWER has served as the primary system supporting a team of six members in overseeing physical health checks for the entire active caseload of over 51,000 patients within the Trust. VIEWER provided a comprehensive overview of physical health checks by leveraging NLP-based information extraction methods. Through interactive visualisations, team members can readily identify patients who require a physical health assessment. Each identified patient is provided with a hyperlink that directs clinicians to their profile within the EHR system, facilitating actions such as scheduling a check or contacting GPs for additional information. After completion, these checks are recorded in the source EHR system and then incorporated into VIEWER. This systematic approach has streamlined the check process and resulted in an increase in completed assessments. For example, Figure 4 shows the rise in the number of measures completed by the team after adopting VIEWER as a systematic approach within a short period of two months.
3.4.2 Medication review